# Meronymic Ontology Extraction via Large Language Models

### v1

@@ Abstract
Ontologies have become essential in today’s digital age as a way of organising the vast amount of readily available unstructured text. In providing formal structure to this information, ontologies have immense value and application across various domains, e.g., e-commerce, where countless product listings necessitate proper product organisation. However, the manual construction of these ontologies is a time-consuming, expensive and laborious process. In this paper, we harness the recent advancements in large language models (LLMs) to develop a fully-automated method of extracting product ontologies, in the form of meronymies, from raw review texts. We demonstrate that the ontologies produced by our method surpass an existing, BERT-based baseline when evaluating using an LLM-as-a-judge. Our investigation provides the groundwork for LLMs to be used more generally in (product or otherwise) ontology extraction.
@@ 1 Introduction
The proliferation of the internet has led to an ever-increasing amount of unstructured text data, presenting both opportunities and challenges in harnessing this wealth of information. Ontologies are one method of organising the information from unstructured text and formally representing it as a graph. Such data representations are important in numerous downstream tasks, such as review aggregation Konjengbam et al. (2018), sentiment analysis Schouten and Frasincar (2018) and product question answering Kulkarni et al. (2019). Often, these ontologies can be as simple as consisting only of part-whole relations, i.e. they are meronymies, which have been shown to be particularly useful in review aggregation contexts concerning product Oksanen et al. (2021), travel Rago et al. (2025) or movie Cocarascu et al. (2019) recommendation. However, the manual construction of (even simple meronymic) ontologies is not only time-consuming but often demands domain expertise.
@@
To this end, research efforts have been directed at learning these ontologies from data, notably via deep learning methods (see Al-Aswadi et al. (2020) for an overview). More recently, large language models (LLMs) Min et al. (2024), which have revolutionised the field of NLP, have also found application in this domain with its increasing adoption in key ontology learning tasks of term and relation extraction Ye et al. (2022). However, most of these efforts have been concentrated on extracting taxonomic (is), rather than meronymic (is part of), relations. While taxonomies are routinely applied in increasingly specific classification of entities, in review aggregation contexts, meronomies are particularly useful for representing features of entities. For example, Oksanen et al. (2021) introduce a method for extracting meronomies using two fine-tuned BERTs Devlin et al. (2019). However, the method still relies on some manual annotation from humans. Further, the effective evaluation of these meronymic ontologies is also non-trivial, since no ground-truth examples exist for this task currently.
@@
Figure 1: Our complete pipeline for extracting meronymic ontologies from product reviews using LLMs, consisting of the four tasks of aspect, synset, concept and relation extraction. In the example shown here, we begin with a review from a customer. The first task comprises the extraction of the relevant aspects from the review, i.e. watch, strap and wristband, with a fine-tuned LLM. The second task then uses a word embedding to extract synset groups, i.e. discarding wristband since it is a synonym of strap. The third task is to select the most relevant of these with an LLM to obtain the concepts. Finally, in the last task, a fine-tuned LLM extracts the relations between concepts, i.e. identifying that strap is part of (in meronymic relation with) watch, resulting in the final product ontology.
@@
In this paper, we harness the recent advancements in LLMs to make a number of contributions improving on the pipeline for ontology extraction proposed by Oksanen et al. (2021). Concretely:
@@
$\bullet$ We empirically evaluate our LLM-based approach against a BERT-based method (Oksanen et al., 2021), finding significant gains in relevance. [^1] Source code: https://anonymous.4open.science/r/llm-ontology-8C78/README.md
@@ 2 Meronymic Ontology Extraction via LLMs
In this section, we first describe the dataset used in both the method and the evaluation, as well as the LLMs used for the different extraction tasks. We then describe our approach to the four tasks constituting the meronymic ontology extraction pipeline, namely: aspect extraction, synset extraction, concept extraction and relation extraction. Our approach is illustrated and motivated in Figure 1.
@@
The product reviews were obtained from the Amazon Reviews 2023 dataset Hou et al. (2024). For the generation of the product ontologies, we selected the same five products as Oksanen et al. (2021) to allow for a fair comparison. These products were chosen to represent the diverse selection of products on Amazon: video games, televisions, necklaces/watches, and stand mixers. We randomly selected 100,000 reviews for each product, except for stand mixers, where only 26,464 reviews were available. The full dataset of 100,000 reviews was used for the synset extraction task. However, for the aspect and relation extraction tasks (which relied on LLMs), we limited the input to only 1,000 reviews due to the processing time of LLMs.
@@
For all our extraction tasks, we used the Mistral-7B-Instruct-v0.2222https://huggingface.co/mistralai/Mistral-7B-Instruct-v0.2 LLM. We chose this model as, despite its modest size, it performs well on various natural language tasks and even outperforms its larger counterparts, namely Llama 1 and Llama 2, on some tasks Jiang et al. (2023). Given this, the model is particularly suitable for our use case where computational resources are limited. In all LLM-related tasks, we constrained the LLMs’ outputs to follow JavaScript Object Notation (JSON) grammar by using the decoding framework of Geng et al. (2023), who demonstrate improved performance from enforcing more structured outputs.
@@
To generate a list of aspects from the review texts, we adopted a fine-tuning approach. In Appendix A.2 we report a comparison with prompt-based approaches, which underperformed. For the fine-tuning, we used the SemEval-2014 Task 4 dataset Pontiki et al. (2014), specifically the manually annotated ground-truth test set from the two provided domains, laptops and restaurants, consisting of 1,600 samples, split evenly. We provide training details in Appendix A.1.
@@
For grouping the aspects into synsets, we used Equidistant Nodes Clustering (ENC) on the cosine similarity between the word embeddings of the aspects, which achieves state-of-the-art performance (Chernoskutov and Ustalov, 2018; Oksanen et al., 2021). We also explored K-Means clustering, which we found produced overly large clusters. We report a comparison of these two methods in Appendix B.2.
To produce the word embeddings, we fine-tuned the FastText model (details in Appendix B.1) due to its resilience to noisy data Bojanowski et al. (2017). This is especially crucial in the context of product reviews, which tend to have a higher proportion of misspelled words and abbreviations. Furthermore, unlike the widely used Word2Vec, which may fail to produce vectors for multi-word phrases with lower occurrence, FastText can construct meaningful representations of these phrases by combining multiple subword units.
@@
The shortlisted synset groups obtained from the previous task form the nodes of the final ontology. To construct the ontology, we extracted the relations between these groups. We did so by isolating sentences containing exactly two aspects from different synset groups. These sentences, along with the two aspects, were then used to prompt the LLM to determine whether a part-whole relationship existed between them (see Appendix D.2 for the prompt). This effectively reduces the task to a multiple-choice problem with answers: 1) the first aspect is a part of the second aspect, 2) the second aspect is a part of the first aspect, and 3) there is no such relationship between the two aspects. We believe this presents a reasonable starting point for meronymies and leave to future work an investigation of more complex relationships between multiple aspects.
We used distillation to fine-tune the Mistral model for this task on 1000 synthetic samples output by Gemini (which was prompted as above) and using the same training settings as those in Appendix A.1. We generated these samples from five product categories (backpack, cardigan, camera, guitar and laptop) that are different from those we tested the model on to ensure it does not simply memorise the relationships.
@@ 3 Experimental Evaluation
We compared against the baseline method proposed by Oksanen et al. (2021). To ensure a fair comparison, we generated two versions of our ontology: an unedited version (full) and a shortened version (short), which has an equal number of terms as the baseline. This was done by keeping the top-K most commonly occurring terms in our ontology, where K is set to the number of terms in the baseline ontology. All experiments were run on an NVIDIA GeForce RTX 4090 GPU. We report the mean results over three runs.
@@
Since user studies are costly for evaluation, we left them to future work and instead turned to LLM-as-a-judge, as is increasingly accepted for simple, domain-specific tasks Huang et al. (2025); Wang et al. (2025). We used Gemini 1.5 Flash [^4] https://ai.google.dev/gemini-api/docs/models#gemini-1.5-flash as the LLM judge, to which we provided a detailed scoring guide that included descriptions of what each score means and examples for each. The LLM judge then used this guide to output a score from 1 to 5. In addition, the LLM judge was prompted to provide explanations for its scoring as part of the chain-of-thought process.
@@ 4 Conclusions
Limitations
The evaluations in this paper rely on LLM-as-a-judge; supplementing this with user studies may provide more evidence for our findings. For example, in the real-world, well-organised product ontologies may ultimately be used to drive user engagement. A comprehensive user study could therefore extend to the impact of product ontologies on downstream metrics. We note, however, that user studies may suffer from a different set of biases (Chen et al., 2024). Product ontologies currently also lack a standard benchmark, with which future progress could be better evaluated. Finally, the performance-improving benefits of LLM-based approaches need to be weighed against potentially higher costs, such as in environmental or monetary terms. As with most LLM-based approaches, hallucinations are a challenge in our work. While a fundamental solution to this problem is outside the scope of this paper, we partially mitigate against this by fine-tuning our models on real product ontologies. Future work could also investigate the use of ensembles of LLMs as a way of mitigating hallucinations in any single model.
@@ Acknowledgments
Rago was funded by The Alan Turing Institute on the UK-Italy Trustworthy AI Visiting Researcher Programme (Award Reference: VRP006). The authors also thank Esmanda Wong for her contributions to this work as part of her Master’s thesis.
@@ A.1 Training Details
We fine-tune the LLM on the 1,600 annotated samples from the SemEval-2014 Task 4 dataset that relate to laptops and restaurants. We set aside 10% of the data for validation. We use the Adam optimizer (Kingma and Ba, 2014) with a learning rate of $10^{-4}$ and a cosine scheduler with a warm-up ratio of 0.1. We use LoRA (Hu et al., 2022) applied to linear projection layers with $r=4$ and $\alpha=16$. We train the model for 3 epochs on effective batch sizes of 16 (using gradient accumulation). The model achieved a final accuracy of 0.9909 on the validation set.

### v2

@@ Abstract
Ontologies have become essential in today’s digital age as a way of organising the vast amount of readily available unstructured text. In providing formal structure to this information, ontologies have immense value and application across various domains, e.g., e-commerce, where countless product listings necessitate proper product organisation. However, the manual construction of these ontologies is a time-consuming, expensive and laborious process. In this paper, we harness the recent advancements in large language models (LLMs) to develop a fully-automated method of extracting product ontologies, in the form of meronomies, from raw review texts. We demonstrate that the ontologies produced by our method surpass an existing, BERT-based baseline when evaluating using an LLM-as-a-judge. Our investigation provides the groundwork for LLMs to be used more generally in (product or otherwise) ontology extraction.
@@ 1 Introduction
The proliferation of the internet has led to an ever-increasing amount of unstructured text data, presenting both opportunities and challenges in harnessing this wealth of information. Ontologies are one method of organising the information from unstructured text and formally representing it as a graph. Such data representations are important in numerous downstream tasks, such as review aggregation Konjengbam_18, sentiment analysis Schouten_18 and product question answering Kulkarni_19. Often, these ontologies can be as simple as consisting only of part-whole relations, i.e. they are meronomies, which have been shown to be particularly useful in review aggregation contexts concerning product Oksanen_21, travel Rago_25 or movie Cocarascu_19 recommendation. However, the manual construction of (even simple meronymic) ontologies is not only time-consuming but often demands domain expertise.
@@
To this end, research efforts have been directed at learning these ontologies from data, notably via deep learning methods (see Al-Aswadi_20 for an overview). More recently, large language models (LLMs) Min_24, which have revolutionised the field of NLP, have also found application in this domain with its increasing adoption in key ontology learning tasks of term and relation extraction Ye_22. However, most of these efforts have been concentrated on extracting taxonomic (is), rather than meronymic (is part of), relations. While taxonomies are routinely applied in increasingly specific classification of entities, in review aggregation contexts, meronomies are particularly useful for representing features of entities. For example, Oksanen_21 introduce a method for extracting meronomies using two fine-tuned BERTs Devlin_19. However, the method still relies on some manual annotation from humans. Further, the effective evaluation of these meronymic ontologies is also non-trivial, since no ground-truth examples exist for this task currently.
@@
Figure 1: Our complete pipeline for extracting meronymic ontologies from product reviews using LLMs, consisting of the four tasks of aspect, synset, concept and relation extraction. In the example shown here, we begin with a review from a customer. The first task comprises the extraction of the relevant aspects from the review, e.g. watch, strap and wristband, with a fine-tuned LLM. The second task then uses a word embedding to extract synset groups, e.g. discarding wristband since it is a synonym of strap. The third task is to select the most relevant of these with an LLM to obtain the concepts. Finally, in the last task, a fine-tuned LLM extracts the relations between concepts, i.e. identifying that strap is part of (in meronymic relation with) watch, resulting in the final product ontology.
@@
In this paper, we harness the recent advancements in LLMs to make a number of contributions improving on the pipeline for ontology extraction proposed by Oksanen_21. Concretely:
@@
$\bullet$ We empirically evaluate our LLM-based approach against a BERT-based method (Oksanen_21), finding significant gains in relevance. [^1] Source code: https://github.com/dkaizhang/llm-meronomy
@@ 2 Meronymic Ontology Extraction via LLMs
In this section, we first describe the dataset used in both the method and the evaluation, as well as the LLMs used for the different extraction tasks. We then outline our approach to the four tasks constituting the meronymic ontology extraction pipeline, namely: aspect extraction, synset extraction, concept extraction and relation extraction. Our approach is illustrated and motivated in Figure 1.
@@
The product reviews were obtained from the Amazon Reviews 2023 dataset Hou_24. For the generation of the product ontologies, we selected the same five products as Oksanen_21 to allow for a fair comparison. These products were chosen to represent the diverse selection of products on Amazon: video games, televisions, necklaces/watches and stand mixers. We randomly selected 100,000 reviews for each product, except for stand mixers, where only 26,464 reviews were available. The full dataset of 100,000 reviews was used for the synset extraction task. However, for the aspect and relation extraction tasks (which relied on LLMs), we limited the input to only 1,000 reviews due to the processing time of LLMs.
@@
For all our extraction tasks, we used the Mistral-7B-Instruct-v0.2 LLM [^2] https://huggingface.co/mistralai/Mistral-7B-Instruct-v0.2. We chose this model as, despite its modest size, it performs well on various natural language tasks and even outperforms its larger counterparts, namely Llama 1 and Llama 2, on some tasks Jiang_23. Given this, the model is particularly suitable for our use case where computational resources are limited. In all LLM-related tasks, we constrained the LLMs’ outputs to follow JavaScript Object Notation (JSON) grammar by using the decoding framework of Geng_23, who demonstrate improved performance from enforcing more structured outputs.
@@
To generate a list of aspects from the review texts, we adopted a fine-tuning approach. In Appendix A.2 we report a comparison with prompt-based approaches, which underperformed. For the fine-tuning, we used the SemEval-2014 Task 4 dataset Pontiki_14, specifically the manually annotated ground-truth test set from the two provided domains, laptops and restaurants, consisting of 1,600 samples, split evenly. We provide training details in Appendix A.1.
@@
For grouping the aspects into synsets, we used Equidistant Nodes Clustering (ENC) on the cosine similarity between the word embeddings of the aspects, which achieves state-of-the-art performance (chernoskutov2018equidistant; Oksanen_21). We also explored K-Means clustering, which we found produced overly large clusters. We report a comparison of these two methods in Appendix B.2.
To produce the word embeddings, we fine-tuned the FastText model (details in Appendix B.1) due to its resilience to noisy data Bojanowski_17. This is especially crucial in the context of product reviews, which tend to have a higher proportion of misspelled words and abbreviations. Furthermore, unlike the widely used Word2Vec, which may fail to produce vectors for multi-word phrases with lower occurrence, FastText can construct meaningful representations of these phrases by combining multiple subword units.
@@
The shortlisted synset groups obtained from the previous task form the nodes of the final ontology. To construct the ontology, we extracted the relations between these groups. We did so by isolating sentences containing exactly two aspects from different synset groups. These sentences, along with the two aspects, were then used to prompt the LLM to determine whether a part-whole relationship existed between them (see Appendix D.2 for the prompt). This effectively reduces the task to a multiple-choice problem with answers: 1) the first aspect is a part of the second aspect; 2) the second aspect is a part of the first aspect; and 3) there is no such relationship between the two aspects. We believe this presents a reasonable starting point for meronomies and leave to future work an investigation of more complex relationships between multiple aspects.
We used distillation to fine-tune the Mistral model for this task on 1000 synthetic samples outputted by Gemini (which was prompted as above) and using the same training settings as those in Appendix A.1. We generated these samples from five product categories (backpack, cardigan, camera, guitar and laptop) that are different from those we tested the model on to ensure it does not simply memorise the relationships.
@@ 3 Experimental Evaluation
We compared against the baseline method proposed by Oksanen_21. To ensure a fair comparison, we generated two versions of our ontology: an unedited version (full) and a shortened version (short), which has an equal number of terms as the baseline. This was done by keeping the top-K most commonly occurring terms in our ontology, where K is set to the number of terms in the baseline ontology. All experiments were run on an NVIDIA GeForce RTX 4090 GPU. We report the mean results over three runs.
@@
Since user studies are costly for evaluation, we left them to future work and instead turned to LLM-as-a-judge, as is increasingly being deployed for simple, domain-specific tasks Huang_25; Wang_25. We used Gemini 1.5 Flash [^4] https://ai.google.dev/gemini-api/docs/models#gemini-1.5-flash as the LLM judge, to which we provided a detailed scoring guide that included descriptions of what each score means and examples for each. The LLM judge then used this guide to output a score from 1 to 5. In addition, the LLM judge was prompted to provide explanations for its scoring as part of the chain-of-thought process.
@@ 4 Conclusions
5 Limitations
The evaluations in this paper rely on LLM-as-a-judge; supplementing this with user studies may provide more evidence for our findings. For example, in the real-world, well-organised product ontologies may ultimately be used to drive user engagement. A comprehensive user study could therefore extend to the impact of product ontologies on downstream metrics. We note, however, that user studies may suffer from a different set of biases (chen2024humans). As with most LLM-based approaches, hallucinations are a challenge in our work. While a fundamental solution to this problem is outside the scope of this paper, we partially mitigate against this by fine-tuning our models on real product ontologies. Future work could also investigate the use of ensembles of LLMs as a way of mitigating hallucinations in any single model. Product ontologies currently also lack a standard benchmark, with which future progress could be better evaluated. Finally, the performance-improving benefits of LLM-based approaches need to be weighed against potentially higher costs, such as in environmental or monetary terms.
@@ Acknowledgments
Rago was partially funded by The Alan Turing Institute on the UK-Italy Trustworthy AI Visiting Researcher Programme (Award Reference: VRP006). The authors thank Esmanda Wong for her contributions to this work as part of her Master’s thesis.
@@ A.1 Training Details
We fine-tune the LLM on the 1,600 annotated samples from the SemEval-2014 Task 4 dataset that relate to laptops and restaurants. We set aside 10% of the data for validation. We use the Adam optimizer (kingma2014adam) with a learning rate of $10^{-4}$ and a cosine scheduler with a warm-up ratio of 0.1. We use LoRA (hu2022lora) applied to linear projection layers with $r=4$ and $\alpha=16$. We train the model for 3 epochs on effective batch sizes of 16 (using gradient accumulation). The model achieved a final accuracy of 0.9909 on the validation set.